OpenStudy (anonymous):

Which ordered pair in the form (a, b) is a solution of this equation? 4a – 5b = 24 A. (0, 6) B. (–5, –2) C. (6, 0) D. (–2, –5)

sammixboo (sammixboo):

\(\large\color{ black}{\large {\bbox[5pt, white ,border:2px solid white ]{ \rm W}}}\)\(\large\color{ black}{\large {\bbox[5pt, white ,border:2px solid white ]{ \rm E}}}\)\(\large\color{ black}{\large {\bbox[5pt, white ,border:2px solid white ]{ \rm L}}}\)\(\large\color{ black}{\large {\bbox[5pt, white ,border:2px solid white ]{ \rm C}}}\)\(\large\color{ black}{\large {\bbox[5pt, white ,border:2px solid white ]{ \rm O}}}\)\(\large\color{ black}{\large {\bbox[5pt, white ,border:2px solid white ]{ \rm M}}}\)\(\large\color{ black}{\large {\bbox[5pt, white ,border:2px solid white ]{ \rm E}}}\) \(\large\color{ black}{\large {\bbox[5pt, white ,border:2px solid white ]{ \rm T}}}\)\(\large\color{ black}{\large {\bbox[5pt, white ,border:2px solid white ]{ \rm O}}}\) \(\large\color{ white}{\large {\bbox[5pt,#00c5ff ,border:2px solid white ]{ \rm O}}}\)\(\large\color{ white}{\large {\bbox[5pt,#00c5ff ,border:2px solid white ]{ \rm P}}}\)\(\large\color{ white}{\large {\bbox[5pt,#00c5ff ,border:2px solid white ]{ \rm E}}}\)\(\large\color{ white}{\large {\bbox[5pt,#00c5ff ,border:2px solid white ]{ \rm N}}}\)\(\large\color{ white}{\large {\bbox[5pt,#92dba2 ,border:2px solid white ]{ \rm S}}}\)\(\large\color{ white}{\large {\bbox[5pt,#92dba2 ,border:2px solid white ]{ \rm T}}}\)\(\large\color{ white}{\large {\bbox[5pt,#92dba2 ,border:2px solid white ]{ \rm U}}}\)\(\large\color{ white}{\large {\bbox[5pt,#92dba2 ,border:2px solid white ]{ \rm D}}}\)\(\large\color{ white}{\large {\bbox[5pt,#92dba2 ,border:2px solid white ]{ \rm Y}}}\) We can plug in the different answer choices into the equation given in the problem :) Let's start with Choice A :) (0, 6) - Since it says (a, b) in the problem, we now that 0 = a and b = 6, so we can plug in our information into this equation given : 4a - 5b = 24 4*0 - 5*6 = 24 Following the order of operations, we must multiply first starting from the left side of the equation, so can you tell me what is 4*0?

OpenStudy (anonymous):

0

sammixboo (sammixboo):

Great! So now we have 0 - 5*6 = 24 Now still following the order of operations, we must multiply, so what is 5*6?

OpenStudy (anonymous):

30

sammixboo (sammixboo):

Good so we have 0 - 30 = 24, and we know that 0 - 30 is 30, but does 30 = 24? Nope! So Choice A is not our answer :( Now for Choice B :D (-5, -5) - Since it says (a, b) in the problem, we know that a = -5 and b = -2, so we can plug in our information into this equation given : 4a - 5b = 24 4*-5 - 5*-5 = 24 Following the order of operations, we must multiply first starting from the left side of the equation, so can you tell me what is 4*-5?

OpenStudy (anonymous):

-20..